# Positive traces on certain $\operatorname{SL}(2)$ Coulomb branches

### v1

@@ 1.5 Positive Traces
Let $T$ be a positive trace on $\mathcal{O}_{c}^{\mathbb{D}_{n}}$, and consider its unique even extension to $\mathcal{O}_{c}^{C_{n}}$ given by the even weight function $w_{0}$ and the analytic formula in (EKRS, 21, Proposition 3.1). We wish to show that $T$ is positive on $\mathcal{O}_{c}^{C_{n}}$. Because $T$ is positive on $\mathcal{O}_{c}^{\mathbb{D}_{n}}$, we must have $T(e_{0}R(k)e_{0}\rho(e_{0}R(k)e_{0}))>0$ for any even polynomial $R$, or equivalently
$\int_{i\mathbb{R}}R(z)\bar{R}(-z)w_{0}(z)|\mathrm{d}z|>0.$ (1.3)
@@
Also, letting $P$ be the polynomial such that $x^{n}y^{n}e_{0}=P(k+\frac{n}{2})e_{0}$, we have
$\int_{i\mathbb{R}}R(z)\bar{R}(-z-n)P(z+\tfrac{n}{2})w_{0}(z)|\mathrm{d}z|>0$ (1.4)
@@
for all even polynomials $R$ by Section 1.5. By Lemma 1.5.1(2), Equation 1.3 is equivalent to $w_{0}(z)$ being nonnegative on $i\mathbb{R}$ because for a sequence of polynomials $S_{i}$ tending to some almost everywhere nonnegative even function $U$ in $L^{1}(\mathbb{R},|w_{0}(z)|)$, the sequence of polynomials $\frac{S_{i}(z)+S_{i}(-z)}{2}$ tends to $U$ in $L^{1}(\mathbb{R},|w_{0}(z)|)$ because $w_{0}$ is even.
@@
Also, letting $S$ be the polynomial such that $R(z)=S(z+\frac{n}{2})$ and noting that $P(z+\frac{n}{2})w_{0}(z)$ is holomorphic between $i\mathbb{R}$ and $\frac{n}{2}+i\mathbb{R}$, Equation 1.4 is equivalent to the statement that
$\int_{\frac{n}{2}+i\mathbb{R}}S(z)\bar{S}(-z)P(z)w_{0}(z-\tfrac{n}{2})|\mathrm{d}z|>0$
@@
for all polynomials $S$ such that $S(z)=S(n-z)$. Again using Lemma 1.5.1(2), it follows that $P(z)w_{0}(z-\frac{n}{2})$ is nonnegative on $i\mathbb{R}$.
@@
The positivity conditions obtained on $P$ and $w_{0}$ are equivalent to those obtained in (EKRS, 21, Proposition 4.20) for type A, so the trace $T$ is positive on $\mathcal{O}_{c}^{\mathbb{D}_{n}}$ if and only if it is positive on $\mathcal{O}_{c}^{C_{n}}$. Thus $T$ is the restriction of a positive trace on $\mathcal{O}_{c}^{C_{n}}$, as desired. ∎
Chapter 2 Pure $\operatorname{SL}(2)/\operatorname{PGL}(2)$ gauge theory
2.1 Description of the algebra
@@
Let $v,w$ be variables satisfying $wv=qvw$. These variables generate the $q$-Weyl algebra. We add $v^{-1}$ to this algebra and call the result $W_{0}$. Elements $\{q^{i}v-q^{-i}{v^{-1}}\}$ generate a multiplicative subset $S$ satisfying both the left and the right Ore condition. We localize at $W_{0}$ at $S$ and call the resulting algebra $W$.
Gaiotto and Teschner GT (24) describe the $K$-theoretic Coulomb branch of pure $\operatorname{SL}(2)$ and $\operatorname{PGL}(2)$ gauge theories using the following algebra. Consider the algebra with generators $v,u_{\pm}$ and relations $u_{\pm v}=q^{\pm 1}vu_{\pm},$
$u_{+}u_{-}=\frac{1}{(v-v^{-1})(qv-(qv)^{-1})},\quad u_{-}u_{+}=\frac{1}{(v-v^{-1})(q^{-1}v-qv^{-1})}.$
Then $\mathcal{A}$ is the subalgebra generated by $v^{k}+v^{-k}$ and $v^{k}u_{+}+v^{-k}u_{-}$ for all integers $k$.
The algebra $\mathcal{A}$ can be embedded into $W$ as follows: $v=v,u_{+}=w(v-v^{-1})^{-1}$, $u_{-}=w^{-1}(v-v^{-1})^{-1}$. Indeed, $u_{\pm}v=q^{\pm 1}vu_{\pm}$,
$u_{+}u_{-}=w(v-v^{-1})^{-1}w^{-1}(v-v^{-1})^{-1}=(qv-q^{-1}v^{-1})^{-1}(v-v^{-1})^{-1},$
@@
$u_{-}u_{+}=w^{-1}(v-v^{-1})^{-1}w(v-v^{-1})^{-1}=(q^{-1}v-qv^{-1})^{-1}(v-v^{-1})^{-1}.$
[…]
@@
The elements $H_{a}$ are defined as $H_{a}=q^{\frac{a}{2}}(v^{a}u_{+}+v^{-a}u_{-})$. Gaiotto and Teschner consider the automorphism $\rho(H_{a})=H_{a-2}$ and $g(H_{a})=H_{a-4}$, but the definition of twisted and positive trace in physics is slightly different: $T(ab)=T(g(b)a)$ instead of $T(ab)=T(bg(a))$ and $T(\rho(a)a)>0$ instead of $T(a\rho(a))>0$. Any twisted trace satisfies $T(a)=T(g(a))$. Hence the $g$-twisted trace in one sense is $g^{-1}$-twisted trace in another. So we consider $g(H_{a})=H_{a+4}$ and $g(H_{a})=H_{a+m}$ for even $m$ more generally.
Note that $g$ and $\rho$ can be lifted to $W$ as follows: $g(v)=v$, $g(w)=q^{\frac{m}{2}}v^{m}w=q^{-\frac{m}{2}}wv^{m}$, $\rho(v)=v^{-1}$, $\rho(w)=q^{\frac{m}{4}}v^{-\frac{m}{2}}w^{-1}=q^{-\frac{m}{4}}w^{-1}v^{-\frac{m}{2}}$.
We assume that $0<q<1$ and $m\geq 0$.
[…]
@@
We will write elements of $\mathcal{A}$ and $W$ as $a=\sum w^{i}f_{i}$, where $f_{i}$ is a rational function in $v$. The function $f_{i}$ may have poles only at ±$\pm q^{k}$ for some $k$. We denote the residue of $f_{i}$ at ±$\pm q^{k}$ by $f_{i}^{(k)}$ or $f_{i}^{(k,-)}$ respectively. We will also use notation the $f_{i}^{(k,+)}=f_{i}^{(k)}$
Proposition 2.1.1.
@@
The algebra $\mathcal{A}$ coincides with the set of elements $a=\sum w^{i}f_{i}$ of $W$ such that all $f_{i}$ have simple poles only possibly at ±$\pm q^{j}$, $j\in\mathbb{Z}$ and the corresponding residues $f_{i}^{(j,\pm)}$ satisfy
  1. 1.
    $f_{-i}(v^{-1})=(-1)^{i}f_{i}(v)$;
  2. 2.
@@
  3. 3.
    For a positive integer $k$ we have $f_{i+k}^{(i,\pm)}=(-1)^{k+1}f_{i-k}^{(i,\pm)}$, where $\pm$ is the same sign on both sides.

### v2

@@ 6 Positive traces
Let $T$ be a positive trace on $\mathcal{O}_{c}^{\mathbb{D}_{n}}$, and consider its unique even extension to $\mathcal{O}_{c}^{C_{n}}$ given by the even weight function $w_{0}$ and the analytic formula in [11, Proposition 3.1]. We wish to show that $T$ is positive on $\mathcal{O}_{c}^{C_{n}}$. Because $T$ is positive on $\mathcal{O}_{c}^{\mathbb{D}_{n}}$, we must have $T(e_{0}R(k)e_{0}\rho(e_{0}R(k)e_{0}))>0$ for any even polynomial $R$, or equivalently
$\int_{{\rm i}\mathbb{R}}R(z)\bar{R}(-z)w_{0}(z)|\mathrm{d}z|>0.$ (6.2)
@@
Also, letting $P$ be the polynomial such that $x^{n}y^{n}e_{0}=P\bigl(k+\frac{n}{2}\bigr)e_{0}$, we have
$\int_{{\rm i}\mathbb{R}}R(z)\bar{R}(-z-n)P\bigl(z+\tfrac{n}{2}\bigr)w_{0}(z)|\mathrm{d}z|>0$ (6.3)
@@
for all even polynomials $R$ by (6). By Lemma 6.1 (2), (6.2) is equivalent to $w_{0}(z)$ being nonnegative on $\rm i$$\mathbb{R}$ because for a sequence of polynomials $S_{i}$ tending to some almost everywhere nonnegative even function $U$ in $L^{1}(\mathbb{R},|w_{0}(z)|)$, the sequence of polynomials $\frac{S_{i}(z)+S_{i}(-z)}{2}$ tends to $U$ in $L^{1}(\mathbb{R},|w_{0}(z)|)$ because $w_{0}$ is even.
@@
Also, letting $S$ be the polynomial such that $R(z)=S(z+\frac{n}{2})$ and noting that $P(z+\frac{n}{2})w_{0}(z)$ is holomorphic between $\rm i$$\mathbb{R}$ and $\frac{n}{2}+{\rm i}\mathbb{R}$, (6.3) is equivalent to the statement that
$\int_{\frac{n}{2}+{\rm i}\mathbb{R}}S(z)\bar{S}(-z)P(z)w_{0}\bigl(z-\tfrac{n}{2}\bigr)|\mathrm{d}z|>0$
@@
for all polynomials $S$ such that S(z)=S(nz)S(z)=S(n-z). Again using Lemma 6.1 (2), it follows that $P(z)w_{0}\bigl(z-\frac{n}{2}\bigr)$} is nonnegative on $\rm i$$\mathbb{R}$.
@@
The positivity conditions obtained on $P$ and $w_{0}$ are equivalent to those obtained in [11, Proposition 4.20] for type A, so the trace $T$ is positive on $\mathcal{O}_{c}^{\mathbb{D}_{n}}$ if and only if it is positive on $\mathcal{O}_{c}^{C_{n}}$. Thus $T$ is the restriction of a positive trace on $\mathcal{O}_{c}^{C_{n}}$, as desired. ∎
Part 2: Pure $\boldsymbol{\operatorname{SL}(2)/\operatorname{PGL}(2)}$ gauge theory
7 Description of the algebra
7.1 Definition of $\boldsymbol{\mathcal{A}}$, connection to the BFN construction
Let $0<q<1$ be a number. Consider the algebra $\mathcal{B}$ with generators $v$, $v^{-1}$, $u_{\pm}$ and relations $u_{\pm}v=q^{\pm 1}vu_{\pm}$},
$u_{+}u_{-}=\frac{1}{\bigl(v-v^{-1}\bigr)\bigl(qv-(qv)^{-1}\bigr)},\qquad u_{-}u_{+}=\frac{1}{\bigl(v-v^{-1}\bigr)\bigl(q^{-1}v-qv^{-1}\bigr)}.$
Definition 7.1.
$\mathcal{A}$ is the subalgebra of $\mathcal{B}$ generated by elements $v^{k}+v^{-k}$ and $H_{a}=q^{\tfrac{a}{2}}(v^{a}u_{+}+v^{-a}u_{-})$ for all positive integers $k$ and all integers $a$.
Gaiotto and Teschner [15] use $\mathcal{A}$ to describe the $K$-theoretic Coulomb branch of pure $\operatorname{SL}(2)$ and $\operatorname{PGL}(2)$ gauge theories. We will explain the connection from a mathematical point of view below.
First, we note that $\mathcal{A}$ can be embedded into a localized $q$-Weyl algebra.
Definition 7.2.
@@ 7.1 Definition of $\boldsymbol{\mathcal{A}}$, connection to the BFN construction
Let $W_{0}=\mathbb{C}\langle v,w\rangle/(wv-qvw)$ be a $q$-Weyl algebra. Then elements $v$, $w$ and $\big\{q^{i}v-q^{-i}{v^{-1}}\big\}$ generate a multiplicative subset $S$ satisfying both the left and the right Ore condition. We define $W$ to be the Ore localization of $W$ at $S$.
Lemma 7.3.
There is an embedding of $\mathcal{B}$ into $W$ given by $v\mapsto v$, $u_{+}\mapsto w\bigl(v-v^{-1}\bigr)^{-1}$ and $u_{-}\mapsto w^{-1}\bigl(v-v^{-1}\bigr)^{-1}$.
Proof.
We see that the images satisfy the defining relations $u_{\pm}v=q^{\pm 1}uv_{\pm}$ of $\mathcal{B}$. For the other two relations, we have
$u_{+}u_{-}=w\bigl(v-v^{-1}\bigr)^{-1}w^{-1}\bigl(v-v^{-1}\bigr)^{-1}=\bigl(qv-q^{-1}v^{-1}\bigr)^{-1}\bigl(v-v^{-1}\bigr)^{-1},$
@@
$u_{-}u_{+}=w^{-1}\bigl(v-v^{-1}\bigr)^{-1}w\bigl(v-v^{-1}\bigr)^{-1}=\bigl(q^{-1}v-qv^{-1}\bigr)^{-1}\bigl(v-v^{-1}\bigr)^{-1}.$
The basis of $\mathcal{B}$ is given by expressions $v^{l}$, $u_{+}^{k}v^{l}$ and $u_{-}^{k}v^{l}$ for positive integers $k$ and integers $v_{l}$. Each $k$ corresponds to its own power of $w$, then different $l$ give linearly independent elements. ∎
We identify $\mathcal{A}$ with its image in $W$. This is the only property of $\mathcal{A}$ that we will use in Sections 8 and 9. We discuss the connection to the BFN construction for completeness.
We turn to BFN Coulomb branches. Pure gauge theory means that $N=0$. In this case, BFN construction gives equivariant $K$-theory of the affine Grassmannian
$\mathcal{A}^{q}_{G,N}=K^{G(\mathbb{C}[[t]])\rtimes\mathbb{C}_{q}^{\times}}(\operatorname{Gr}_{G}).$
The group $\mathbb{C}_{q}^{\times}$ acts by loop rotation on $G(O)$ and on $\operatorname{Gr}_{G}$.
Let $q$ be a formal parameter instead of a number, the definition of $\mathcal{A}$ works with the following change: $\mathcal{A}$ is a $\mathbb{C}\big[q,q^{-1}\big]$-algebra and we change the generator $H_{a}$ to $\tilde{H}_{a}=(v^{a}u_{+}+v^{-a}u_{-})$. Then $\mathcal{A}$ should be equal to $K^{\operatorname{SL}(2,\mathbb{C}[[t]])\rtimes\mathbb{C}_{q}^{\times}}(\operatorname{Gr}_{\operatorname{PGL}_{2}})$, a modification of BFN construction. We use this as a motivation to consider $\mathcal{A}$, but a rigorous proof is beyond the scope of this paper.
{lemma*}
We have $\mathcal{A}\cong K^{\operatorname{SL}(2,\mathbb{C}[[t]])\rtimes\mathbb{C}_{q}^{\times}}(\operatorname{Gr}_{\operatorname{PGL}_{2}})$. More precisely, we identify $W$ with $K^{\mathbb{C}^{\times}_{v}\times\mathbb{C}^{\times}_{q}}\allowbreak\times(\operatorname{Gr}_{T})_{\rm loc}$, the localized quantized Coulomb branch of pure $\mathbb{C}^{\times}_{v}$ gauge theory. Then the image of the localization map
$(i_{*})^{-1}\colon\ K^{\operatorname{SL}(2,\mathbb{C}[[t]])\rtimes\mathbb{C}_{q}^{\times}}(\operatorname{Gr}_{\operatorname{PGL}_{2}})\to K^{\mathbb{C}^{\times}_{v}\times\mathbb{C}^{\times}_{q}}(\operatorname{Gr}_{T})_{\rm loc}$
equals to the image of $\mathcal{A}$ in $W$ under the map $u_{+}\mapsto\bigl(v-v^{-1}\bigr)^{-1}w$, $u_{-}\mapsto\bigl(v-v^{-1}\bigr)^{-1}w^{-1}$.
Sketch of the proof.
Note that we switched between left and right fractions compared to Lemma 7.3, but the proof of the inclusion $\mathcal{B}\hookrightarrow W$ works without change.
Similarly to the proof of [4, Lemma 6.9], one can prove that $K^{\operatorname{SL}(2,\mathbb{C}[[t]])\rtimes\mathbb{C}_{q}^{\times}}(\operatorname{Gr}_{\operatorname{PGL}_{2}})$ is generated by the dressed miniscule monopole operators and the K-theory of a point. Then we use the localization map, it is also called the abelianization map in physics literature. For pure gauge theory this map is constructed in [3, Section 6.3]. We get the map
$(i_{*})^{-1}\colon\ K^{\operatorname{SL}(2,\mathbb{C}[[t]])\rtimes\mathbb{C}_{q}^{\times}}(\operatorname{Gr}_{\operatorname{PGL}_{2}})\to K^{\mathbb{C}^{\times}_{v}\times\mathbb{C}^{\times}_{q}}(\operatorname{Gr}_{T})_{\rm loc},$
where $\mathbb{C}^{\times}_{v}\subset\operatorname{SL}(2,\mathbb{C})$ is the maximal torus and the multiplicative subset on the right is defined so that $W=K^{\mathbb{C}^{\times}_{v}\times\mathbb{C}^{\times}_{q}}(\operatorname{Gr}_{T})_{\rm loc}$. Let $v$ be the generator of the $\mathbb{C}^{\times}_{v}$-equivariant $K$-theory of a point corresponding to the standard representation, similarly for $q$ and $\mathbb{C}^{\times}_{q}$. Then the image of $K^{\operatorname{SL}(2)}(pt)$ under the localization map consists of Laurent polynomials in $v$ symmetric with respect to $v\mapsto v^{-1}$. Dressed miniscule monopole operators are defined as $K$-theory classes of the sheaves $O(l)$, $l\in\mathbb{Z}$ on the smooth $\operatorname{PGL}_{2}(\mathbb{C}[[t]])$-orbit $\mathbb{P}^{1}\subset\operatorname{Gr}_{\operatorname{PGL}_{2}}$. Here the $\operatorname{SL}_{2}$-equivariant structure is the natural one, and the action of $\mathbb{C}_{q}^{\times}$ on $O(l)$ is trivial. It can be checked that the image of $[O(l)]$ in $W$ is
$\bigl(v-v^{-1}\bigr)^{-1}\bigl(v^{l+1}w-v^{-l-1}w^{-1}\bigr)=-\tilde{H}_{l+1}.$
This finishes the proof. ∎
We claim that $K^{\operatorname{SL}(2,\mathbb{C}[[t]])\rtimes\mathbb{C}_{q}^{\times}}(\operatorname{Gr}_{\operatorname{PGL}_{2}})$ contains Coulomb branches of pure $\operatorname{SL}(2)$ and $\operatorname{PGL}(2)$ gauge theories. The inclusion
$\mathcal{A}_{\operatorname{SL}_{2},0}^{q}=K^{\operatorname{SL}(2,\mathbb{C}[[t]])\rtimes\mathbb{C}_{q}^{\times}}(\operatorname{Gr}_{\operatorname{SL}_{2}})\subset K^{\operatorname{SL}(2,\mathbb{C}[[t]])\rtimes\mathbb{C}_{q}^{\times}}(\operatorname{Gr}_{\operatorname{PGL}_{2}})$
comes from the inclusion of the affine Grassmannians. Let $\tau_{1}\colon\mathcal{A}\to\mathcal{A}$ be the map that fixes $v$ and sends $w$ to w-w. Then $\mathcal{A}_{\operatorname{SL}_{2},0}^{q}$ is the subalgebra $\mathcal{A}^{\tau_{1}}$ of $\tau_{1}$-fixed elements. The inclusion
$\mathcal{A}_{\operatorname{PGL}_{2},0}^{q}=K^{\operatorname{PGL}(2,\mathbb{C}[[t]])\rtimes\mathbb{C}_{q}^{\times}}(\operatorname{Gr}_{\operatorname{PGL}_{2}})\subset K^{\operatorname{SL}(2,\mathbb{C}[[t]])\rtimes\mathbb{C}_{q}^{\times}}(\operatorname{Gr}_{\operatorname{PGL}_{2}})$
follows from the fact that any $\operatorname{PGL}_{2}$-equivariant sheaf is $\operatorname{SL}_{2}$-equivariant via the projection $\operatorname{SL}_{2}\twoheadrightarrow\operatorname{PGL}_{2}$. In terms of the generators above, $\mathcal{A}_{\operatorname{PGL}_{2},0}^{q}$ is generated by $v^{n}+v^{-n}$ with $n$ even and $H_{a}$ with $a$ odd. Any such element is fixed by the involution $\tau_{2}$ such that $\tau_{2}(v)=-v$, $\tau_{2}(w)=-w$ and we have $\mathcal{A}^{q}_{\operatorname{PGL}_{2},0}=\mathcal{A}^{\tau_{2}}$.
@@
Gaiotto and Teschner consider the automorphism $\rho(H_{a})=H_{a-2}$ and $g(H_{a})=H_{a-4}$, but the definition of twisted and positive trace in physics is slightly different: $T(ab)=T(g(b)a)$ instead of $T(ab)=T(bg(a))$ and $T(\rho(a)a)>0$ instead of $T(a\rho(a))>0$. Any twisted trace satisfies T(a)=T(g(a)){T(a)=T(g(a))}. Hence, a $g$-twisted trace in one sense is a $g^{-1}$-twisted trace in another. So we consider $g(H_{a})=H_{a+4}$ and, more generally, $g(H_{a})=H_{a+m}$ and $\rho(H_{a})=H_{a+\tfrac{m}{2}}$ for even $m$.
We classify twisted and positive traces on $\mathcal{A}$ in the case when $m\geq 0$.
For future use, we note that $g$ and $\rho$ can be extended to an (anti-)automorphism of $W$ as follows:
$g(v)=v,\qquad$ $g(w)=q^{\frac{m}{2}}v^{m}w=q^{-\frac{m}{2}}wv^{m},$
$\rho(v)=v^{-1},\qquad$ $\rho(w)=q^{\frac{m}{4}}v^{-\frac{m}{2}}w^{-1}=q^{-\frac{m}{4}}w^{-1}v^{-\frac{m}{2}}.$
Remark 7.4.
There is an isomorphism between $\mathcal{A}_{q}$ and $\mathcal{A}_{q^{-1}}$ that does not change $w$ and sends $v$ to $v^{-1}$. This isomorphism interchanges $g_{m}$-twisted traces for $\mathcal{A}_{q}$ and $g_{-m}$-twisted traces for $\mathcal{A}_{q^{-1}}$. It can be shown that in the integral formula from Proposition 8.2 we can take $q$ to be a formal parameter and take formal power series $\omega(z)$ instead of an actual quasi-periodic function. This would give a classification of twisted traces for all $q$ and hence for all $m$. For 0<q<1{0<q<1} and negative $m$, most of these power series will have quadratic exponential growth and converge nowhere, similarly to the series for Jacobi theta function $\sum q^{n^{2}}z^{n}$ with $q>1$ instead of $0<q<1$. Because of this and the fact that $m<0$ case is less physically significant, we work only with the case $m\geq 0$.
Remark 7.5.
Geometrically, $g_{m}$ should correspond to the twist by $m$-th power of the determinant bundle on the affine Grassmannian.
7.2 Description of the image $\boldsymbol{\mathcal{A}\subset W}$
@@
We will write elements of $\mathcal{A}$ and $W$ as $a=\sum w^{i}f_{i}=\sum_{i=-N}^{N}w^{i}f_{i}$, where $N$ is a nonnegative integer and $f_{i}$ is a rational function in $v$. The function $f_{i}$ may have poles only at $\pm q^{k}$ for some $k$. We denote the residue of $f_{i}$ at $\pm q^{k}$ by $f_{i}^{(k)}$ or $f_{i}^{(k,-)}$, respectively. We will also use notation the $f_{i}^{(k,+)}=f_{i}^{(k)}$
Proposition 7.6.
@@
The algebra $\mathcal{A}$ coincides with the set of elements $a=\sum w^{i}f_{i}$ of $W$ such that all $f_{i}$ have simple poles only possibly at $\pm q^{j}$, $j\in\mathbb{Z}$ and the corresponding residues $f_{i}^{(j,\pm)}$ satisfy
  1. $(1)$
    $f_{-i}\bigl(v^{-1}\bigr)=(-1)^{i}f_{i}(v)$;
  2. $(2)$
@@
  3. $(3)$
    for a positive integer $k$, we have $f_{i+k}^{(i,\pm)}=(-1)^{k+1}f_{i-k}^{(i,\pm)}$, where $\pm$ is the same sign on both sides.